# On the Optimization Dynamics of RLVR: Gradient Gap and Step Size Thresholds

### v2

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR), which uses simple binary feedback to post-train large language models, has shown significant empirical success. However, a principled understanding of why it works has been lacking. This paper builds a theoretical foundation for RLVR by analyzing its training process at both the full-response (trajectory) and token levels. Central to our analysis is a quantity called the Gradient Gap, which formalizes the direction of improvement from low-reward to high-reward regions of the response space. We prove that convergence critically depends on aligning the update direction with this Gradient Gap. Moreover, we derive a sharp step-size threshold based on the magnitude of the Gradient Gap: below it, learning converges, whereas above it, performance collapses. Our theory further predicts how the critical step size must scale with response length and the success rate, thereby explaining why practical heuristics such as length normalization improve stability and showing that, with a fixed learning rate, the success rate can stagnate strictly below $100\%$. We validate these predictions through controlled bandit simulations.
@@ 1 Introduction
Large language models (LLMs) have recently achieved significant advances through reinforcement learning post-training, which aligns them with complex tasks and preferences (ziegler2019fine, ; ouyang2022, ; shao2024deepseekmath, ; team2025kimi, ). In particular, Reinforcement Learning with Verifiable Rewards (RLVR) has emerged as a powerful approach for post-training LLMs on tasks where success can be automatically checked (e.g. using a compiler or solver) (guo2025deepseek, ). RLVR methods have shown impressive empirical gains by leveraging binary success/failure feedback instead of human judgments, thereby simplifying the RL pipeline. Techniques in this vein (e.g. variants of Proximal Policy Optimization, a.k.a. PPO (schulman2017proximal, ), like GRPO (shao2024deepseekmath, ), DAPO (yu2025dapo, ), Dr. GRPO (liu2025understanding, ), etc.) eliminate the need for learned reward or value models, relying purely on verifiable outcome signals. This has enabled LLMs to achieve state-of-the-art results on challenging reasoning and code-generation tasks, demonstrating the promise of RLVR-driven fine-tuning.
However, empirical progress in RLVR has far outpaced our theoretical understanding. The optimization process remains largely a black box: we do not fully understand why RL-based post-training works so well, under what conditions it might falter, or how to tune it for stable convergence. Recent PPO-based variants (e.g. GRPO, DAPO, Dr. GRPO) have sprung up to improve training stability, each introducing different heuristics like normalizing gradient updates by the output length or standardizing rewards by the group’s variance. Yet it remains unclear which of these design choices truly matter; without a principled basis, their adoption is guided more by intuition than by theory. This gap is especially pronounced given RLVR’s sparse binary rewards (each episode yields just a single success/failure bit), which make it difficult to analyze how gradient descent navigates the model’s vast parameter space or how the policy’s output distribution shifts toward higher-reward answers. In practice, practitioners often resort to trial-and-error for critical hyperparameters and algorithmic choices, where a mis-tuning can destabilize training or even cause catastrophic collapse (e.g., forgetting pre-trained knowledge or converging to trivial outputs). These challenges underscore the need for a rigorous theoretical framework to demystify RLVR’s optimization dynamics and reduce reliance on guesswork.
@@
  • Unified RLVR theory: We develop a principled framework for RLVR under binary rewards, introducing the Gradient Gap to characterize the improvement direction from low- to high-reward responses.
@@ 1.1 Related Works
Recent efforts in RL-based language model post-training have introduced a family of GRPO-style algorithms that extend or modify Proximal Policy Optimization (PPO) for verifiable feedback settings. GRPO itself eliminates the value critic by estimating advantages from a group of sampled responses, using relative reward normalization instead of a learned baseline (shao2024deepseekmath, ). Building on this idea, DAPO augmented GRPO by decoupling the PPO clipping range and dynamically filtering out cases where all responses in a batch are correct or all are incorrect (yu2025dapo, ). Dr. GRPO revisits the advantage normalization procedure, arguing that removing length and variance normalizations (i.e. using only a mean baseline) can prevent bias in policy updates (liu2025understanding, ). Additional related papers are discussed in Appendix˜A.
In parallel, theoretical work has established convergence guarantees for policy gradient (PG) methods, including REINFORCE and actor-critic algorithms. In finite Markov decision processes with softmax policies, the PG objective often satisfies a PL condition, implying that any stationary point is globally optimal and that vanilla gradient ascent converges at a sublinear rate (agarwal2021theory, ; xiao2022convergence, ). Actor-critic methods also achieve provable convergence by using two-timescale updates or pessimistic value estimation (wu2020finite, ; zanette2021provable, ). However, extending these guarantees to post-training large language models with verifiable binary rewards remains challenging, as sparse success/failure signals provide very limited gradient information.
@@ 2 Problem Set-Up
The model generates tokens in an autoregressive fashion: at every step $t$, the next token $o_{t}$ is sampled conditioned on the prompt $q$ and all previously generated tokens $\vec{\boldsymbol{o}}\!\,_{<t}=(o_{1},o_{2},\ldots,o_{t-1})$. Formally, $\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,\mid q)=\prod_{t=1}^{|\vec{\boldsymbol{o}}\!\,|}\uppi_{\theta}(o_{t}\mid q,\vec{\boldsymbol{o}}\!\,_{<t})$. Each conditional distribution is defined by a softmax over token logits $\boldsymbol{h}_{\theta}(\cdot)$:
@@
${\rm maximize}_{\theta\in\mathds{R}^{d}}\quad J(\uppi_{\theta})\;:\,=\;\mathds{E}_{q\sim\mathds{P}(Q),\,\vec{\boldsymbol{o}}\!\,\sim\uppi_{\theta}(\cdot\mid q)}\big[r^{\star}(q,\vec{\boldsymbol{o}}\!\,)\big]\,.$ (2)
@@
This generic formulation captures a broad family of post-training algorithms used in RLVR. Representative examples are:
@@
The classical policy gradient method updates parameters in the direction
$\nabla_{\theta}\,J(\uppi_{k})\;=\;\mathds{E}_{q\sim\mathds{P}(Q),\,\vec{\boldsymbol{o}}\!\,\sim\uppi_{k}(\cdot\mid q)}\big[A(q,\vec{\boldsymbol{o}}\!\,)\cdot\nabla_{\theta}\,\log\uppi_{k}(\vec{\boldsymbol{o}}\!\,\mid q)\big]\,$, (4)
where the advantage function is given by $A(q,\vec{\boldsymbol{o}}\!\,)=r^{\star}(q,\vec{\boldsymbol{o}}\!\,)-\mathds{E}_{\vec{\boldsymbol{o}}\!\,^{\prime}\sim\uppi_{k}(\cdot\mid q)}[r^{\star}(q,\vec{\boldsymbol{o}}\!\,^{\prime})]$. In this case, the update rule $\theta_{t+1}=\theta_{k}+\alpha\cdot\nabla_{\theta}\,J(\uppi_{k})$ can be rewritten in our generic form by setting $\boldsymbol{w}_{k}=\nabla_{\theta}\,J(\uppi_{k})/\|\nabla_{\theta}\,J(\uppi_{k})\|_{2}$ and $\eta_{k}=\alpha\,\|\nabla_{\theta}\,J(\uppi_{k})\|_{2}$. Viewed in this way, Dr. GRPO (liu2025understanding, ) emerges as a variant that replaces the single-sample advantage with a group-wise demeaned version.
@@
GRPO has recently become a standard choice for RLVR. The full algorithm incorporates clipping ratios and multi-step updates (see Section˜B.1). To connect it with the generic policy gradient form, we consider a simplified one-step approximation without clipping. In this case, the gradient direction is
$\boldsymbol{g}_{\mathrm{GRPO}}(\uppi_{k})\;=\;\mathds{E}_{q\sim\mathds{P}(Q),\,\vec{\boldsymbol{o}}\!\,\sim\uppi_{k}(\cdot\mid q)}\Big[\frac{A(q,\vec{\boldsymbol{o}}\!\,)}{\sigma(q)}\cdot\frac{1}{|\vec{\boldsymbol{o}}\!\,|}\nabla_{\theta}\,\log\uppi_{k}(\vec{\boldsymbol{o}}\!\,\mid q)\Big]\,$, (5)
@@ 3 Trajectory-Level Analysis
In this section, we study the optimization scheme (3) on a single prompt $q$. We take a trajectory-level view, where each response $\vec{\boldsymbol{o}}\!\,$ is treated as a single unit rather than a sequence of tokens. By abstracting away the internal structure, the analysis becomes simpler yet still revealing. We begin by outlining the key ingredients of this trajectory-level view, then examine both its success modes and failure cases. Although this setup is only a warm-up for the more detailed token-level analysis, it already highlights several nontrivial and illuminating properties of RLVR.
@@ 3.1 Key Ingredients: Gradient Gap and Gap Alignment
Recall that the optimization objective is the correction rate of the model $\uppi_{\theta}$ on prompt $q$: $J_{q}(\uppi_{\theta}):\,=\mathds{E}_{\vec{\boldsymbol{o}}\!\,\sim\uppi_{\theta}(\cdot\mid q)}\big[r^{\star}(q,\vec{\boldsymbol{o}}\!\,)\bigm|q\big]$. To analyze this, we partition the response space $\mathcal{O}$ into two sets based on the verifiable reward $r^{\star}(q,\cdot)$:
$\mathcal{O}_{q}^{+}\;:\,=\;\big\{\vec{\boldsymbol{o}}\!\,\in\mathcal{O}\bigm|r^{\star}(q,\vec{\boldsymbol{o}}\!\,)=1\big\}\qquad\mbox{and}\qquad\mathcal{O}_{q}^{-}\;:\,=\;\big\{\vec{\boldsymbol{o}}\!\,\in\mathcal{O}\bigm|r^{\star}(q,\vec{\boldsymbol{o}}\!\,)=0\big\}\,$, (6)
@@
$\boldsymbol{g}_{q}^{+}(\uppi_{\theta}):\,=\mathds{E}_{\vec{\boldsymbol{o}}\!\,\sim\uppi_{\theta}^{+}(\cdot\mid q)}\big[\nabla_{\theta}\,\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,\mid q)\big]~~~\mbox{and}~~~\boldsymbol{g}_{q}^{-}(\uppi_{\theta}):\,=\mathds{E}_{\vec{\boldsymbol{o}}\!\,\sim\uppi_{\theta}^{-}(\cdot\mid q)}\big[\nabla_{\theta}\,\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,\mid q)\big].$ (8)
@@
Crucially, the Gradient Gap is directly proportional [^1] A formal proof of this is found in Section B.2 to the policy gradient given in equation (4):
@@
This shows that the Gradient Gap captures the true direction of improvement. Unlike the full policy gradient $\nabla_{\theta}\,J_{q}(\uppi_{\theta})$, $\boldsymbol{g}_{q}^{+}(\uppi_{\theta})-\boldsymbol{g}_{q}^{-}(\uppi_{\theta})$ is not scaled down by the variability factor $J_{q}(1-J_{q})$, making it a purer indicator of where to move.
@@
If $\|\boldsymbol{w}_{t}\|_{2}=1$, this equals $\Delta\mu_{q}(k)\;=\;\|\boldsymbol{g}_{q}^{+}(k)-\boldsymbol{g}_{q}^{-}(k)\|_{2}\cdot\cos\angle\big\{\boldsymbol{w}_{t},\boldsymbol{g}_{q}^{+}(k)-\boldsymbol{g}_{q}^{-}(k)\big\}$, which depends both on the magnitude of the Gradient Gap and the angle of alignment.
@@ 3.2 Main Findings
We now turn to the central findings of our analysis. Proofs will be deferred to Appendices˜C and D. Before presenting the results, let us impose a mild regularity condition on the policy score function.
@@
    $\big\|\nabla_{\theta}\,\log\uppi_{\theta^{\prime}}(\vec{\boldsymbol{o}}\!\,\mid q)-\nabla_{\theta}\,\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,\mid q)\big\|_{2}\;\leq\;L_{\mathrm{o}}\cdot\|\theta^{\prime}-\theta\|_{2}.$ (13)
Throughout this section, we use the shorthand $J_{q}(k)=J_{q}(\uppi_{k})$ to denote the performance at iteration $k$.

### v3

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR), which uses simple binary feedback to post-train large language models, has found significant empirical success. However, a principled understanding of why it works is lacking. This paper builds a theoretical foundation for RLVR by analyzing its training process at both the full-response (trajectory) and token levels. Central to our analysis is a new quantity called the Gradient Gap, which formalizes the direction of improvement from low-reward to high-reward regions of the response space. We prove that convergence critically depends on aligning the update direction with this Gradient Gap. Moreover, we derive a sharp step-size threshold based on the magnitude of the Gradient Gap: below it, learning converges, whereas above it, performance collapses. Our theory further predicts how the critical step size must scale with response length and the success rate, thereby explaining why practical heuristics such as length normalization improve stability and showing that, with a fixed learning rate, the success rate can stagnate strictly below $100\%$. Importantly, our theory holds flexibly for any policy-gradient algorithm and so characterizes the dynamics of popular approaches such as REINFORCE and GRPO. We validate these predictions through controlled bandit simulations and language model experiments on post-training Qwen2.5-Math-7B with GRPO.
On the Optimization Dynamics of RLVR:
Gradient Gap and Step Size Thresholds
Joe Suk Yaqi Duan
j.suk@nyu.edu yd2878@stern.nyu.edu
Department of Technology, Operations and Statistics
Stern School of Business, New York University
@@ 1 Introduction
Large language models (LLMs) have recently achieved significant advances through reinforcement learning post-training, which aligns them with complex tasks and preferences (Ziegler et al., 2019; Ouyang et al., 2022; Shao et al., 2024; Team et al., 2025). In particular, Reinforcement Learning with Verifiable Rewards (RLVR) has emerged as a powerful approach for post-training LLMs on tasks where success can be automatically checked (Guo et al., 2025). RLVR methods have shown impressive empirical gains by leveraging binary success/failure feedback instead of human judgments, thereby simplifying the RL pipeline. Techniques in this vein (e.g. variants of Proximal Policy Optimization, a.k.a. PPO (Schulman et al., 2017), like GRPO (Shao et al., 2024), DAPO (Yu et al., 2025), Dr. GRPO (Liu et al., 2025), etc.) eliminate the need for learned reward or value models, relying purely on verifiable outcome signals. This has enabled LLMs to achieve state-of-the-art results on challenging reasoning and code-generation tasks, demonstrating the promise of RLVR-driven fine-tuning.
However, empirical progress in RLVR has far outpaced our theoretical understanding. The optimization process remains largely a black box: we do not fully understand why RL-based post-training works so well, under what conditions it might falter, or how to tune it for stable convergence. Recent PPO-based variants (e.g. GRPO, DAPO, Dr. GRPO) have sprung up to improve training stability, each introducing different heuristics like normalizing gradient updates by the output length or standardizing rewards by the group’s variance. Yet it remains unclear which of these design choices truly matter; without a principled basis, their adoption is guided more by intuition than by theory. This gap is especially pronounced given RLVR’s sparse binary rewards (each episode yields just a single success/failure bit), which make it difficult to analyze how gradient descent navigates the model’s vast parameter space or how the policy’s output distribution shifts toward higher-reward answers. In practice, practitioners often resort to trial-and-error for critical hyperparameters and algorithmic choices, where a mis-tuning can destabilize training or even cause catastrophic collapse (e.g., forgetting pre-trained knowledge or converging to trivial outputs). These challenges underscore the need for a rigorous theoretical framework to demystify RLVR’s optimization dynamics.
@@
  • Unified RLVR theory: We develop a principled framework for RLVR under binary rewards, introducing the Gradient Gap to characterize the improvement direction from low- to high-reward responses. Our theory applies to any policy gradient-based method, and thus describes training behavior of methods such as REINFORCE and GRPO.
@@
  • Empirical validation: We validate our theory through bandit simulations and LLM experiments, including fine-tuning Qwen2.5-Math-7B on GSM8K, DeepScaleR, and DAPO-17k datasets with GRPO, demonstrating close alignment between theory and practice.
@@ 1.1 Related Works
Recent efforts in RL-based language model post-training have introduced a family of GRPO-style algorithms that extend or modify Proximal Policy Optimization (PPO) for verifiable feedback settings. GRPO itself eliminates the value critic by estimating advantages from a group of sampled responses, using relative reward normalization instead of a learned baseline (Shao et al., 2024). Building on this idea, DAPO augmented GRPO by decoupling the PPO clipping range and dynamically filtering out cases where all responses in a batch are correct or all are incorrect (Yu et al., 2025). Dr. GRPO revisits the advantage normalization procedure, arguing that removing length and variance normalizations (i.e. using only a mean baseline) can prevent bias in policy updates (Liu et al., 2025). Additional related papers are discussed in Appendix A.
In parallel, theoretical work has established convergence guarantees for policy gradient (PG) methods, including REINFORCE and actor-critic algorithms. In finite Markov decision processes with softmax policies, the PG objective often satisfies a PL condition, implying that any stationary point is globally optimal and that vanilla gradient ascent converges at a sublinear rate (Agarwal et al., 2021; Xiao, 2022). Actor-critic methods also achieve provable convergence by using two-timescale updates or pessimistic value estimation (Wu et al., 2020; Zanette et al., 2021). However, extending these guarantees to post-training large language models with verifiable binary rewards remains challenging, as sparse success/failure signals provide very limited gradient information.
Theoretical Analyses of Post-Training.
Prior works on theoretical convergence guarantees of post-training focus on upper bounding the norm of reward gradients (Ge et al., 2025; Pang and Jin, 2025; Huang et al., 2025), which for non-convex objectives is a weaker notion of convergence than direct rates on the reward improvement, as we show in this work. Furthermore, these works are constrained to a one-step trajectory-level analysis and thus do not take into account response length normalization in rates. In other directions, Mroueh (2025) study the fixed-point solutions of GRPO and Yuan et al. (2025) conduct a trajectory-level convergence analysis of value-based methods.
@@ 2 Problem Set-Up
The model generates tokens in an autoregressive fashion: at every step $t$, the next token $o_{t}$ is sampled conditioned on the prompt $q$ and all previously generated tokens $\vec{\boldsymbol{o}}\!\,_{<t}=(o_{1},o_{2},\ldots,o_{t-1})$. Formally, $\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,\mid q)=\prod_{t=1}^{|\vec{\boldsymbol{o}}\!\,|}\uppi_{\theta}(o_{t}\mid q,\vec{\boldsymbol{o}}\!\,_{<t})$. Each conditional distribution is defined by a softmax over token logits $\boldsymbol{h}_{\theta}(\cdot)$: [^1] Our convergence results require only mild regularity conditions, and not the softmax parametrization. We present this standard form as the motivating example.
@@
$\max_{\theta\in\mathds{R}^{d}}\quad J(\uppi_{\theta})\;:\,=\;\mathds{E}_{q\sim\mathds{P}_{Q},\,\vec{\boldsymbol{o}}\!\,\sim\uppi_{\theta}(\cdot\mid q)}\big[r^{\star}(q,\vec{\boldsymbol{o}}\!\,)\big]\,.$ (2)
@@
This generic formulation captures a broad family of post-training RVLR algorithms. Representative examples are:
@@
The classical policy gradient method updates parameters in the direction $\nabla_{\theta}\,J(\uppi_{k})=$
$\mathds{E}_{\begin{subarray}{l}q\sim\mathds{P}_{Q},\\ \vec{\boldsymbol{o}}\!\,\sim\uppi_{k}(\cdot\mid q)\end{subarray}}\big[A(q,\vec{\boldsymbol{o}}\!\,)\nabla_{\theta}\,\log\uppi_{k}(\vec{\boldsymbol{o}}\!\,\mid q)\big]\,$, (4)
where the advantage function is given by $A(q,\vec{\boldsymbol{o}}\!\,)=r^{\star}(q,\vec{\boldsymbol{o}}\!\,)-\mathds{E}_{\vec{\boldsymbol{o}}\!\,^{\prime}\sim\uppi_{k}(\cdot\mid q)}[r^{\star}(q,\vec{\boldsymbol{o}}\!\,^{\prime})]$. In this case, the update rule $\theta_{k+1}=\theta_{k}+\alpha\cdot\nabla_{\theta}\,J(\uppi_{k})$ can be rewritten in our generic form by setting $\boldsymbol{w}_{k}=\nabla_{\theta}\,J(\uppi_{k})/\|\nabla_{\theta}\,J(\uppi_{k})\|_{2}$ and $\eta_{k}=\alpha\,\|\nabla_{\theta}\,J(\uppi_{k})\|_{2}$. Viewed in this way, Dr. GRPO (Liu et al., 2025) emerges as a variant that replaces the single-sample advantage with a group-wise demeaned version.
@@
GRPO has recently become a standard choice for RLVR. The full algorithm incorporates clipping ratios and multi-step updates (see Subsection B.1). To connect it with the generic policy gradient form, we consider a simplified one-step approximation without clipping. In this case, the gradient direction is $\boldsymbol{g}_{\mathrm{GRPO}}(\uppi_{k})\;=\;$
$\mathds{E}_{\begin{subarray}{l}q\sim\mathds{P}_{Q},\\ \vec{\boldsymbol{o}}\!\,\sim\uppi_{k}(\cdot\mid q)\end{subarray}}\Big[\frac{A(q,\vec{\boldsymbol{o}}\!\,)}{\sigma(q)}\cdot\frac{1}{|\vec{\boldsymbol{o}}\!\,|}\nabla_{\theta}\,\log\uppi_{k}(\vec{\boldsymbol{o}}\!\,\mid q)\Big]\,$, (5)
@@ 3 Trajectory-Level Analysis
In this section, we study the optimization scheme (3) on a single prompt $q$, which we later generalize to the multiple prompt setting in Section 5. As a warmup, we take a trajectory-level view, where each response $\vec{\boldsymbol{o}}\!\,$ is treated as a single unit rather than a sequence of tokens. By abstracting away the internal structure, the analysis becomes simpler yet still revealing. We begin by outlining the key ingredients of this trajectory-level view, then examine both its success modes and failure cases. Although this setup is only a warm-up for the more detailed token-level analysis, it already highlights several nontrivial and illuminating properties of RLVR.
@@ 3.1 Key Ingredients: Gradient Gap and Gap Alignment
Our optimization objective is the correction rate of the model $\uppi_{\theta}$ on prompt $q$: $J_{q}(\uppi_{\theta}):\,=\displaystyle\mathop{\mathds{E}}_{\vec{\boldsymbol{o}}\!\,\sim\uppi_{\theta}(\cdot\mid q)}\big[r^{\star}(q,\vec{\boldsymbol{o}}\!\,)\bigm|q\big]$. To analyze this, we partition the response space $\mathcal{O}$ into two sets based on the verifiable reward $r^{\star}(q,\cdot)$:
$\mathcal{O}_{q}^{+}$ $\;:\,=\;\big\{\vec{\boldsymbol{o}}\!\,\in\mathcal{O}\bigm|r^{\star}(q,\vec{\boldsymbol{o}}\!\,)=1\big\}$
$\mathcal{O}_{q}^{-}$ $\;:\,=\;\big\{\vec{\boldsymbol{o}}\!\,\in\mathcal{O}\bigm|r^{\star}(q,\vec{\boldsymbol{o}}\!\,)=0\big\}\,.$
@@
$\boldsymbol{g}_{q}^{+}(\uppi_{\theta})$ $:\,=\mathds{E}_{\vec{\boldsymbol{o}}\!\,\sim\uppi_{\theta}^{+}(\cdot\mid q)}\big[\nabla_{\theta}\,\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,\mid q)\big]$
$\boldsymbol{g}_{q}^{-}(\uppi_{\theta})$ $:\,=\mathds{E}_{\vec{\boldsymbol{o}}\!\,\sim\uppi_{\theta}^{-}(\cdot\mid q)}\big[\nabla_{\theta}\,\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,\mid q)\big].$ (7)
@@
Crucially, the Gradient Gap is directly proportional [^2] A formal proof of this is found in Subsection B.2 to the policy gradient given in equation (4):
@@
This shows the Gradient Gap captures the true direction of improvement and, unlike the policy gradient $\nabla_{\theta}\,J_{q}(\uppi_{\theta})$, is not scaled down by the variability factor $J_{q}(1-J_{q})$, making it a purer indicator of where to move.
@@
If $\|\boldsymbol{w}_{k}\|_{2}=1$, this equals $\Delta\mu_{q}(k)\;=\;\|\boldsymbol{g}_{q}^{+}(k)-\boldsymbol{g}_{q}^{-}(k)\|_{2}\cdot\cos\angle\big\{\boldsymbol{w}_{k},\boldsymbol{g}_{q}^{+}(k)-\boldsymbol{g}_{q}^{-}(k)\big\}$, which depends both on the magnitude of the Gradient Gap and the angle of alignment.
@@ 3.2 Main Findings
We now turn to the central findings of our analysis. Proofs will be deferred to Appendices C and F. We first impose mild regularity assumptions on the policy score function, standard in theoretical analyses of policy gradient methods (Agarwal et al., 2021) and also appearing in recent theoretical works on post-training (Huang et al., 2025; Pang and Jin, 2025; Foster et al., 2025; Chen et al., 2025a).
@@
    $\hskip-20.00003pt\big\|\nabla_{\theta}\,\log\uppi_{\theta^{\prime}}(\vec{\boldsymbol{o}}\!\,\mid q)-\nabla_{\theta}\,\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,\mid q)\big\|_{2}\;\leq\;L_{\mathrm{o}}\|\theta^{\prime}-\theta\|_{2}.$
For language models with transformer architectures, Assumption 1 holds on any compact parameter space due to the smoothness of softmax policies and activation functions. In practice, weight decay and gradient clipping ensure fine-tuning operates within a bounded region of parameter space.
Throughout this section, we use shorthand $J_{q}(k)=J_{q}(\uppi_{k})$ to denote the performance at iteration $k$.